OpenStudy (anonymous):

44% of US adults 18-34 years think that "Made In America" ads boost sales. Another poll show that 78% of adults in the 18-34 years category use social media online. Suppose that 85% of US adults in the 18-34 years category think that "Made in America" ads boost sales or use social media online. If US adult in the 18-34 years category is randomly selected A: What is the probability that the person thinks that " Made in America" ads boost sales and uses social media online? B: What is the probability that made in America ads boost sales given that the person does not use social media onl

OpenStudy (anonymous):

I came up with answer p(nus)= .44+.78=11.3

OpenStudy (slaaibak):

p(A or B) = P(A) + P(B) - P(A and B) so P(A and B) = P(A) + P(B) - P(A or B) = .44 + .78 - .85 = 0.37 b) P(A|notB) = P(A and not B) / P(notB)

OpenStudy (anonymous):

OK thanks

OpenStudy (anonymous):

Next, according to reports by Scarborough research, the average monthly house hold cellular phone bill is 73.00 with. A SD of 11.35. What is the probability randomly the bill is more than 100.? I came up with 100-77/2=27%

OpenStudy (anonymous):

I meant 100-73/2 or 100-73/11.35

OpenStudy (anonymous):

Next, according to reports by Scarborough research, the average monthly house hold cellular phone bill is 73.00 with. A SD of 11.35. What is the probability randomly the bill is more than 100.? I came up with 100-77/2=27%